OpenStudy (anonymous):

are m and 2m like terms

OpenStudy (anonymous):

Yes they are.

OpenStudy (anonymous):

yes. If you are adding them they become 3m

OpenStudy (anonymous):

nahh..... haha... In multi step equations....

OpenStudy (saifoo.khan):

m + 2m = (1 +2)m = 3m

OpenStudy (anonymous):

heres the problem... m-4=2m im solving multi step

OpenStudy (anonymous):

would you make m 1m?

OpenStudy (radar):

yes and then m=-4

OpenStudy (anonymous):

yes, if they have the same variable they are. so m +2m=3m

OpenStudy (anonymous):

its a binomial equation means having 2 terms